Find two consecutive negative integers whose product is 90
-thanks

Found 2 solutions by nerdybill, Alan3354:
Answer by nerdybill(7384)

Find two consecutive negative integers whose product is 90
.
Let x = first negative answer
then
x+1 = second negative answer
x(x+1) = 90
x^2 + x = 90
x^2 + x - 90 = 0
(x+10)(x-9) = 0
x = (-10, 9)
Since we're looking for a negative answer, throw out the positive answer leaving:
x = -10 (first negative number)
.
Second negative number"
x+1 = -10+1 = -9
.
Answer: -10 and -9

Answer by Alan3354(69443)

The square root of 90 is apx 9.5, the center
--> -10 and -9.